OpenStudy (anonymous):

Use the equation v=10/p to determine the pressure when the volume is 12 liters.

OpenStudy (anonymous):

if you are to strictly follow that equation then the answer is pretty simple! idk if they want your volume to be converted or not if not then just plug in the values! p= 10/12 =5/6